How many grams of CO2 are produced when 100 g C4H10 undergoes combustion with 200 g O2?

Use the following molar masses: C4H10=58g/mol, O2=32 g/mol and CO2=44g/mol.


1
Expert's answer
2021-04-19T05:09:26-0400

2(C4H10) + 13O2 → 8CO2 + 10H2O


Moles of "C_4H_{10 }= \frac{100}{58}=1.72moles"


Moles of "O_2= \frac{200}{32}=6.25moles"


Moles of "CO_2= \frac{1.72×8}{2}=6.88moles"


Mass of "CO_2= 6.88\u00d744=302.72g"